# Differentially Private Data Generation with Missing Data

### v2

@@ 6.1. Experimental Setup
Datasets. We run our experiments on four tabular datasets as described in Table 1: (i) Adult dataset (Dua and Graff, 2017), which contains information about 32561 individuals from the 1994 US Census (ii) Bank dataset (Moro et al., 2014), which has 45211 rows about direct marketing campaigns of a Portuguese banking institution (iii) BR2000 (Zhang et al., 2014), which consists of 38,000 census records collected from Brazil in the year 2000, (iv) National dataset (Task et al., 2022) from NIST Diverse Community Excerpts which contains information about 15012 individuals US census data. Each dataset has a combination of numerical and categorical columns which are pre-processed according to the synthetic data generation algorithm as discussed in their respective research paper – numerical attributes are discretized into 10 uniform bins or scaled between 0 to 1 and the categorical attributes are encoded using one-hot or ordinal encoding.
Baselines. We consider several existing differentially private data generation methods that do not consider missing data: PrivBayes (Zhang et al., 2014) and AIM (McKenna et al., 2022) from statistical approaches, DPCTGAN (Fang et al., 2022) and DPautoGAN (Tantipongpipat et al., 2021) from deep learning techniques, and Kamino (Ge et al., 2021), which is a mixed approach. These methods have been published in well-known conferences with their code readily available. For each method, we construct the following baselines to deal with missing data as discussed in Section 4.1. The first baseline referred by its original name combines the original method with the complete row-only approach. Next, we construct baselines for the imputation first approach, which first imputes the data and then generates the synthetic data. We initially considered differentially private data imputation using k-means (Clifton et al., 2022) and OLS regression (Das et al., 2022), but they fail in working with categorical columns and require a large privacy budget for the imputation process itself. Hence, we adopt the following imputation methods: (i) random imputation, which fills missing values randomly from the attribute domain; (ii) mean imputation, replacing missing numerical attribute values with the mean and sampling from the probability distribution for categorical attributes; and (iii) Kamino imputation, which leverages intermediate models from Kamino for imputation and shares similar model as Holoclean (Wu et al., 2020; hol, line). Random impute requires no privacy budget, but the other two require splitting of the privacy budget for imputation and for the data generation algorithm.
@@
Parameters. The utility of the algorithms replies upon multiple hyperparameters, and we try our best to tune these parameters by running grid searches. Although we don’t consider privacy costs for tuning, it’s crucial to tune these parameters privately in practice (Mohapatra et al., 2021; Papernot and Steinke, [n.d.]). Deep learning methods typically require a larger privacy budget for meaningful results. Hence, we assign PrivBayes and Kamino $\epsilon=1$, and GAN approaches $\epsilon=3$ by default. Adaptive methods maintain the same budget as their non-adaptive counterparts. PrivBayes operates under pure DP ($\delta=0$), while others employ approximate DP. The $\delta$ value is approximately set one magnitude lower than $1/|D|$ to the nearest exponent of 10.
@@
Missing data. We implement a pipeline that can generate different categories of missing data at different missing percentages, using the approach from Muzellec et al. (Muzellec et al., 2020). MCAR missing data is generated by masking the original dataset using a realization of a Bernoulli variable with a fixed parameter such that there is exactly the required number of missing values. To generate MAR data, we first use 50% of the attributes in the dataset as features for a logistic regression model. The other attributes then have missing values according to random weights in the logistic model. A bias term is fitted using line search to attain the desired proportion of missing values. The MNAR approach works similarly to MAR, with the difference that the 50% non-missing attributes are masked by an MCAR mechanism. This imposes the logistic model to depend potentially on missing values, hence enforcing MNAR missingness. As 50% of the columns in the MAR mechanism has complete values, it comparatively has more complete rows in comparison to MCAR and MNAR. We run experiments on all kinds of missing types for every dataset and go up to 30% missing values except the national dataset, where we stop at 20% due to the lower number of rows in the original dataset.
@@
Metrics. We use two utility metrics for evaluation. The first metric is $k$-way marginal distance. We calculate all $k$-sized combinations of attributes in the dataset (Zhang et al., 2014; Ge et al., 2021), and then report the average variational distance between the true vs the synthetic marginals. For each $k$-sized combination set of attributes in $A\in\mathcal{A}$, we calculate the marginal, $h:\mathcal{D}\rightarrow\mathbb{R}^{|\mathcal{D}(A)|}$ and report the average variational distance between as $\max_{a\in\mathcal{D}(A)}|h(D^{\prime})[a]-h(\bar{D})[a]|$ where $D^{\prime}$ and $\bar{D}$ are the synthetic data and ground true data respectively (Zhang et al., 2014; Ge et al., 2021). In our evaluation, we set $k=1,2$. Smaller values of this metric show more closeness between the true and synthetic data. Our second metric is model training. We consider 9 classification models (LogisticRegression, AdaBoost, GradientBoost, XGBoost, RandomForest, BernoulliNB, DecisionTree, Bagging, and MLP) to classify each attribute in the dataset (e.g., income is more than 50k or not, the loan should be given or not to user) using all other attributes as features. Each target attribute is processed to be a binary attribute, and we try to balance the two classes as much as possible. The quality of the learning task is represented by the average of all models across all attributes. The F1 score is reported for learning quality. Each model is trained using 70% of the synthetic data, and the F1 score is evaluated using 30% of the true data (Ge et al., 2021). Larger values of this metric show better performance. For all our experiments, we run 3 times and report the mean with standard deviation.
@@ 6.2.1. Benchmark Existing DP Methods on Missing Data
[…]
@@ 6.2.3. Evaluate Adaptive Recourse Approach
Varying privacy budget. We show the impact of the privacy budget in Figure 9 for adaptive methods with 10% MCAR missing data on Adult. We vary the privacy budget $\epsilon\in[0.5,1,3,5,10,\infty]$ where $\epsilon=\infty$ refers to the non-private run. First, we note that increasing the privacy budget improves the utility of the synthetic dataset across all algorithms. Second, we observe that PrivBayesE and KaminoI outperformed GAN-based approaches at all privacy budgets. Similar observations are found in Figure 6 and Figure 7. The poor performance of the GAN-based approaches is primarily due to the requirement for specific preprocessing steps and extensive hyperparameter tuning, which was challenging. Third, we find that there is no clear winner for all datasets, all utility metrics, and all privacy budgets, which is consistent with the observations from prior benchmarks on DP synthetic data generation (Tao et al., 2021). However, we observe that PrivBayesE performs better at smaller epsilon values ($\epsilon\leq 1$), but KaminoI has better performance with higher epsilon ($\epsilon>1$). We attribute this phenomenon to the fact that KaminoI trains multiple intermediate models, and these models generally require a higher privacy budget. Therefore, we recommend PrivBayesE at a low privacy regime ($\epsilon\leq 1$) and KaminoI at a higher privacy regime ($\epsilon>1)$.
@@ 7. Related Work
Differentially private synthetic data generation has been studied vastly in prior literature (Bowen and Liu, 2020; Fan, 2020; Zhu et al., 2017; of Standards and Technology, 2018). Prior works generate synthetic data via statistical approaches which estimate low-dimensional marginal distributions (Qardaji et al., 2014; Xiao et al., 2011), deep learning approaches (Frigerio et al., 2019; Jordon et al., 2019), or the combination of the two (Ge et al., 2021). However, all of these algorithms focus on the no missing data setting. Some prior work look into missing data imputation for private datasets but are either not in the differential privacy setting (Patki et al., 2016; Huang et al., 2018; Jagannathan and Wright, 2008) or do not support generating synthetic data as a part of their work (Clifton et al., 2022; Krishnan et al., 2016; Das et al., 2022). Patki et al. (Patki et al., 2016) propose the synthetic data vault framework, which identifies and repairs inconsistencies in the generated synthetic data from incomplete data but does not consider any privacy guarantees. Their approach uses low-way marginals and learns them using Gaussian copulas that may be enhanced using our partial marginal observation approach if made private. Huang et al. (Huang et al., 2018) and Jagannathan et al. (Jagannathan and Wright, 2008) imputation of missing data as a cleaning algorithm for private datasets but consider privacy definitions of k-anonymity and cryptographic distributive computing, respectively. Other works focus on privately imputing missing values. The existing differentially private solutions (Krishnan et al., 2016; Clifton et al., 2022; Das et al., 2022) focus on data imputation and are difficult to adapt for the synthetic data generation problem. PrivateClean (Krishnan et al., 2016) requires a human in the loop that can specify the user-defined specific imputation functions to clean the database. The k-means based word by Clifton et al. (Clifton et al., 2022) and imputation first approach in Das et al. (Das et al., 2022) use part of the privacy budget to impute the missing values using OLS regression. We note that such OLS regression and k-means models cannot be applied to categorical attributes, and using them for imputation incurs a significant amount of privacy budget and leaves little budget for data synthesis.

### v3

@@ 6.1. Experimental Setup
Datasets. We run our experiments on four tabular datasets as described in Table 1: (i) Adult dataset (uci_repo), which contains information about 32561 individuals from the 1994 US Census (ii) Bank dataset (bank), which has 45211 rows about direct marketing campaigns of a Portuguese banking institution (iii) BR2000 (DBLP:conf/sigmod/ZhangCPSX14), which consists of 38,000 census records collected from Brazil in the year 2000, (iv) National dataset (national) from NIST Diverse Community Excerpts which contains information about 15012 individuals US census data. Each dataset has a combination of numerical and categorical columns which are pre-processed according to the synthetic data generation algorithm as discussed in their respective research paper – numerical attributes are discretized into 10 uniform bins or scaled between 0 to 1 and the categorical attributes are encoded using one-hot or ordinal encoding.
Baselines. We consider several existing differentially private data generation methods that do not consider missing data: PrivBayes (DBLP:conf/sigmod/ZhangCPSX14) and AIM (mckenna2022aim) from statistical approaches, DPCTGAN (fang2022dp) and DPautoGAN (tantipongpipat2021differentially) from deep learning techniques, and Kamino (ge2021kamino), which is a mixed approach. These methods have been published in well-known conferences with their code readily available. For each method, we construct the following baselines to deal with missing data as discussed in Section 4.1. The first baseline referred by its original name combines the original method with the complete row-only approach. Next, we construct baselines for the imputation first approach, which first imputes the data and then generates the synthetic data. We initially considered differentially private data imputation using k-means (DBLP:journals/tissec/CliftonHMM22) and OLS regression (DBLP:journals/corr/abs-2206-15063), but they fail in working with categorical columns and require a large privacy budget for the imputation process itself. Hence, we adopt the following imputation methods: (i) random imputation, which fills missing values randomly from the attribute domain; (ii) mean imputation, replacing missing numerical attribute values with the mean and sampling from the probability distribution for categorical attributes; and (iii) Kamino imputation, which leverages intermediate models from Kamino for imputation and shares similar model as Holoclean (aimnet; holocleancode). Random impute requires no privacy budget, but the other two require splitting of the privacy budget for imputation and for the data generation algorithm.
@@
Parameters. The utility of the algorithms replies upon multiple hyperparameters, and we try our best to tune these parameters by running grid searches. Although we don’t consider privacy costs for tuning, it’s crucial to tune these parameters privately in practice (mohapatra2021role; papernothyperparameter). Deep learning methods typically require a larger privacy budget for meaningful results. Hence, we assign PrivBayes and Kamino $\epsilon=1$, and GAN approaches $\epsilon=3$ by default. Adaptive methods maintain the same budget as their non-adaptive counterparts. PrivBayes operates under pure DP ($\delta=0$), while others employ approximate DP. The $\delta$ value is approximately set one magnitude lower than $1/|D|$ to the nearest exponent of 10.
@@
Missing data. We implement a pipeline that can generate different categories of missing data at different missing percentages, using the approach from Muzellec et al. (muzellec2020missing). MCAR missing data is generated by masking the original dataset using a realization of a Bernoulli variable with a fixed parameter such that there is exactly the required number of missing values. To generate MAR data, we first use 50% of the attributes in the dataset as features for a logistic regression model. The other attributes then have missing values according to random weights in the logistic model. A bias term is fitted using line search to attain the desired proportion of missing values. The MNAR approach works similarly to MAR, with the difference that the 50% non-missing attributes are masked by an MCAR mechanism. This imposes the logistic model to depend potentially on missing values, hence enforcing MNAR missingness. As 50% of the columns in the MAR mechanism has complete values, it comparatively has more complete rows in comparison to MCAR and MNAR. We run experiments on all kinds of missing types for every dataset and go up to 30% missing values except the national dataset, where we stop at 20% due to the lower number of rows in the original dataset.
@@
Metrics. We use two utility metrics for evaluation. The first metric is $k$-way marginal distance. We calculate all $k$-sized combinations of attributes in the dataset (DBLP:conf/sigmod/ZhangCPSX14; ge2021kamino), and then report the average variational distance between the true vs the synthetic marginals. For each $k$-sized combination set of attributes in $A\in\mathcal{A}$, we calculate the marginal, $h:\mathcal{D}\rightarrow\mathbb{R}^{|\mathcal{D}(A)|}$ and report the average variational distance between as $\max_{a\in\mathcal{D}(A)}|h(D^{\prime})[a]-h(\bar{D})[a]|$ where $D^{\prime}$ and $\bar{D}$ are the synthetic data and ground true data respectively (DBLP:conf/sigmod/ZhangCPSX14; ge2021kamino). In our evaluation, we set $k=1,2$. Smaller values of this metric show more closeness between the true and synthetic data. Our second metric is model training. We consider 9 classification models (LogisticRegression, AdaBoost, GradientBoost, XGBoost, RandomForest, BernoulliNB, DecisionTree, Bagging, and MLP) to classify each attribute in the dataset (e.g., income is more than 50k or not, the loan should be given or not to user) using all other attributes as features. Each target attribute is processed to be a binary attribute, and we try to balance the two classes as much as possible. The quality of the learning task is represented by the average of all models across all attributes. The F1 score is reported for learning quality. Each model is trained using 70% of the synthetic data, and the F1 score is evaluated using 30% of the true data (ge2021kamino). Larger values of this metric show better performance. For all our experiments, we run 3 times and report the mean with standard deviation.
@@ 6.2.1. Benchmark Existing DP Methods on Missing Data
We demonstrate the impact of missing values on DP synthetic data generation algorithms. Figure 4 shows the performance of the baselines on the Adult dataset with varying levels of missing completely at random (MCAR) data (x-axis). Our results indicate that missing data up to 5% minimally affects the quality of the synthetic data with 1-way and 2-way metrics experiencing 5-19% impact, and F1-score experiencing 1-13% impact. However, as missing data increases, the quality of synthetic data generated by all algorithms degrades drastically. With 20% missing data, the 1-way metric is affected by 14-190%, the 2-way by 18-147%, and the F1-score is impacted by 6-28%. With high amounts of missing data ($\geq 10\%)$, PrivBayes and DPCTGAN exhibit the most decrease in utility. However, Kamino and DPAutoGAN show a consistent level of performance even with more missing data indicating a high degree of stability in the output. The stability of these methods can be attributed to the fact that these methods tend to extract more information from the available data. Kamino learns the functional dependencies of the dataset and tries to preserve them while generating the synthetic data and DPAutoGAN trains an autoencoder as a preprocessing step to learn the low-dimensional statistics of the data.
@@ 6.2.3. Evaluate Adaptive Recourse Approach
Varying privacy budget. We show the impact of the privacy budget in Figure 9 for adaptive methods with 10% MCAR missing data on Adult. We vary the privacy budget $\epsilon\in[0.5,1,3,5,10,\infty]$ where $\epsilon=\infty$ refers to the non-private run. First, we note that increasing the privacy budget improves the utility of the synthetic dataset across all algorithms. Second, we observe that PrivBayesE and KaminoI outperformed GAN-based approaches at all privacy budgets. Similar observations are found in Figure 6 and Figure 7. The poor performance of the GAN-based approaches is primarily due to the requirement for specific preprocessing steps and extensive hyperparameter tuning, which was challenging. Third, we find that there is no clear winner for all datasets, all utility metrics, and all privacy budgets, which is consistent with the observations from prior benchmarks on DP synthetic data generation (tao2021benchmarking). However, we observe that PrivBayesE performs better at smaller epsilon values ($\epsilon\leq 1$), but KaminoI has better performance with higher epsilon ($\epsilon>1$). We attribute this phenomenon to the fact that KaminoI trains multiple intermediate models, and these models generally require a higher privacy budget. Therefore, we recommend PrivBayesE at a low privacy regime ($\epsilon\leq 1$) and KaminoI at a higher privacy regime ($\epsilon>1)$.
@@ 7. Related Work
Differentially private synthetic data generation has been studied vastly in prior literature (Bowen_2020; fan_2020; DBLP:journals/tkde/ZhuLZY17; nist). Prior works generate synthetic data via statistical approaches which estimate low-dimensional marginal distributions (DBLP:conf/sigmod/QardajiYL14; DBLP:journals/tkde/XiaoWG11), deep learning approaches (DBLP:conf/sec/FrigerioOGD19; DBLP:conf/iclr/JordonYS19a), or the combination of the two (ge2021kamino). However, all of these algorithms focus on the no missing data setting. Some prior work look into missing data imputation for private datasets but are either not in the differential privacy setting (patki; huang2018pacas; jagannathan2008privacy) or do not support generating synthetic data as a part of their work (DBLP:journals/tissec/CliftonHMM22; krishnan2016privateclean; DBLP:journals/corr/abs-2206-15063). Patki et al. (patki) propose the synthetic data vault framework, which identifies and repairs inconsistencies in the generated synthetic data from incomplete data but does not consider any privacy guarantees. Their approach uses low-way marginals and learns them using Gaussian copulas that may be enhanced using our partial marginal observation approach if made private. Huang et al. (huang2018pacas) and Jagannathan et al. (jagannathan2008privacy) imputation of missing data as a cleaning algorithm for private datasets but consider privacy definitions of k-anonymity and cryptographic distributive computing, respectively. Other works focus on privately imputing missing values. The existing differentially private solutions (krishnan2016privateclean; DBLP:journals/tissec/CliftonHMM22; DBLP:journals/corr/abs-2206-15063) focus on data imputation and are difficult to adapt for the synthetic data generation problem. PrivateClean (krishnan2016privateclean) requires a human in the loop that can specify the user-defined specific imputation functions to clean the database. The k-means based word by Clifton et al. (DBLP:journals/tissec/CliftonHMM22) and imputation first approach in Das et al. (DBLP:journals/corr/abs-2206-15063) use part of the privacy budget to impute the missing values using OLS regression. We note that such OLS regression and k-means models cannot be applied to categorical attributes, and using them for imputation incurs a significant amount of privacy budget and leaves little budget for data synthesis.